3. (a) Use the Riemann sum to calculate the area under the curve y = −x 2 + 3x + 2 between x = −1 and x = 2. 


b.(b) Let R = \lim _{n\to \infty \:}\left(\sum _{j=1}^{n-1}\:\left(3+\frac{4j}{n}\right)^2.\frac{4}{n}\right), write R as a definite integral and hence evaluate R


c.) If x^y\:=\:e^{x-y}, prove that \frac{d}{dx}=\frac{ln\:x}{\left(1+lnx\right)^2}


1
Expert's answer
2021-05-28T05:26:19-0400

Part a.

"\u222babf(x)dx\u2248\u0394x(f(x0)+f(x1)+f(x2)+\u22ef+f(xn\u22122)+f(xn\u22121))"

"where Δx=\frac{b−a}{n}"

We have that "a=\u22121, b=2, n=4."

"Therefore, Δx=\frac{2−(−1)}{4}=3/4"

Divide the interval "[\u22121,2]"  into n=4 subintervals of the length Δx=3/4 with the following endpoints "a=\u22121 , \u22121\/4, 1\/2, 5\/4, 2=b"

Now, just evaluate the function at the left endpoints of the subintervals.

"f(x0)=f(\u22121)=0"

"f(x1)=f(\u22121\/4)=21\/16=1.3125"

"f(x2)=f(1\/2)=15\/4=3.75"

"f(x3)=f(5\/4)=117\/16=7.3125"

Finally, just sum up the above values and multiply by "\u0394x=3\/4"

"3\/4(0+1.3125+3.75+7.3125)=9.28125"


Part b

"\lim _{n\to \infty \:}\left(\sum _{i=1}^n\left(3+\frac{4n}{n}\right)^2\frac{4}{n}\right)"

"\mathrm{The\:definite\:integral\:is\:defined\:as:\:}\lim _{n\to \infty }\left(\sum _{i=1}^n\left(f\left(c_i\right)\Delta \:x\right)\right)=\int _a^bf\left(x\right)dx"

"\mathrm{Where\:}Δx=\frac{b-a}{n},\:\mathrm{and}\:c_i=a+Δx⋅i"

"=\int _0^1196dx"

"=196"


Part c

"=e^{x-y}\frac{d}{dx}\left(x-y\right)"

"=1-\frac{d}{dx}\left(y\right)"

"=e^{x-y}\left(1-\frac{d}{dx}\left(y\right)\right)"

"\left(\frac{\ln \left(x\right)}{\left(1+\ln \left(x\right)\right)^2}\right)"